OpenStudy (cenaida):

Solve the system, using substitution. 3a - b = 11 and 2a + 3b = 0 A. (3, 2) B. (3, -2) C. (-3, 2) D. (-3, -2) E. no solution

OpenStudy (john_es):

Using the second, \[a=-3b/2\]Plug into the second, \[3(-3b/2)-b=11\Rightarrow -9b/2-b=11\Rightarrow -11b/2=11\Rightarrow b=-2\]Using again, the first equation, \[a=3\]

OpenStudy (cenaida):

thaaaaank yo

OpenStudy (cenaida):

u